The perimeter of a rectangle is 80 feet. The length of the rectangle is 2 feet less than 5 times the width. What is the length and width of the rectangle?
Answer by quentin252(19)

1. length + width = 80
2. length = 5*width - 2
1. length = 80 - width
2. 80 - width = 5*width - 2
-->> 80 = 6*width - 2
-->> 82 = 6*width
-->> 82/6 = 41/3 = width
-->> length = 80 - 41/3 = 240/3 - 41/3 = 199/3
note this = 5*(41/3) - 2 = 205/3-6/3 = 199/3